### v2

@@ 5.2 Main Results
We compare CodeLutra with baselines in code generation for the data query task, as shown in Table 1. We found that existing open-source LLMs like Llama-3-8B still have a significant performance gap in code generation for data queries compared to closed-source LLMs like GPT-4. Although supervised fine-tuning can help bridge this gap—e.g., SFT increases the EX of Llama-3-8B on Spider from 59.3% to 67.9%—there remains a notable difference with GPT-4’s 74.4%. Through our refinement framework, Llama-3-8B after four iterations exceeded SFT performance by 16.9% and even outperforms GPT-4 with an execution accuracy of 76.6%. Additionally, on the more challenging BIRD dataset, after three iterations, CodeLutra significantly improved the EX of the base model from 22.3 to 43.1, achieving performance very close to GPT-4.
@@ 5.3 More experiments
Our framework CodeLutra leverages both positive and negative answer pairs to iteratively improve model performance, particularly by minimizing the generation of incorrect responses. But what happens when we omit the negative samples and rely solely on supervised fine-tuning using positive samples generated by the model? In this ablation, we compare the performance of our objective 2 with a model trained with $\mathcal{L}_{\text{SFT}}(\pi;\mathcal{D}_{t}^{c})$. As seen in Figure 3(a), without negative samples (purple line), the model’s performance plateaus across iterations, remaining close to the baseline. In contrast, incorporating negative samples (blue line) leads to steady performance improvements over successive iterations. This ablation confirms that including negative samples is critical to refining the model’s ability to distinguish between optimal and suboptimal responses, significantly boosting overall performance.
@@
Table 2: Comparison with DPO (fine-tuned on SFT model)
. Methods Spider DS1000 $\mathcal{L}_{\text{DPO}}$ 17.2 12.4 Ours 76.6 48.6
@@ 6 Further Analysis on CodeLutra
Based on the responses of the initial model $\pi_{0}$, we divide the test set into a correct set and an error set. We track the quality trends of the model in these two sets across iterations. Using the cosine similarity metric based on the BLEURT embedding proposed by Sellam et al. (2020), we calculate similarities denoted as $\text{sim}{\pi_{t}}(\hat{y},y_{\text{gt}})$ for the model fine-tuned over $t$ iterations. Here $\hat{y}$ denotes the model generation, and $y_{\text{gt}}$ is the ground truth solution. As shown in Figure 3(b), we observe that the similarity between the model’s output on the correct set and the ground truth remains stable (see purple bars), while with each iteration, the similarity between the model’s output on the error set and the ground truth increases significantly—from 0.48 to 0.54. This indicates that CodeLutra helps the base model improve outputs on error set, while the outputs on correct cases remain qualitatively stable.
@@ 7 Conclusion
Appendix A Extended Literature Discussion
While previous approaches Le et al. (2022); Shojaee et al. (2023); Ni et al. (2024); Chen et al. (2024) leverage execution results predominantly to refine incorrect code until a correct solution emerges, they often rely heavily on complex reward signals from critic models Le et al. (2022); Shojaee et al. (2023) or sophisticated prompt engineering  Chen et al. (2024); Ni et al. (2024). In contrast, our method focuses on using these execution outcomes to probabilistically differentiate correct from incorrect generations via a Bradley–Terry model. This conceptual shift allows our framework to remain simpler, more generalizable, and less dependent on proprietary large-scale models: rather than crafting intricate instructions tailored to advanced architectures or relying on external ground truths like unit tests, we directly let the model learn from failures by contrasting them with successes, thereby improving its coding proficiency over iterative training cycles. Compared other methods, which introduce complexities that may not scale down to smaller models or open-source environments, our minimalistic strategy consistently demonstrates strong performance.
Furthermore, although theoretical insights into dual loss objectives have been discussed in other work Liu et al. (2024), our approach refines and applies these principles to the uniquely challenging domain of code generation—showing that the integration of execution feedback into a direct, probabilistic learning framework can yield practical gains without the pitfalls of overly elaborate interventions. While prior iterative reasoning methods excel at mathematical tasks Pang et al. (2024), code generation poses distinct challenges due to strict syntactic and semantic correctness requirements. In mathematical reasoning, a correct final output does not necessarily guarantee an entirely accurate answer Wang et al. (2024a). However, in code generation, even minor errors in the code can result in completely incorrect execution outcomes. This makes leveraging execution results to construct preference pairs particularly meaningful. Our proposed framework, CodeLutra, addresses these challenges by integrating both successful and failed attempts using a Bradley–Terry model, enabling the refinement of outputs without relying on explicit feedback or ground-truth labels. By generalizing from errors and iteratively improving attempts, CodeLutra consistently reduces syntactic and logical flaws while maintaining the quality of correct solutions. Empirically, it matches or even surpasses advanced models like GPT-4, advancing the democratization of high-quality code generation in a domain with strict correctness requirements.
Input : Training set $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$; Initial base model $\pi_{0}$; Number of code responses per input $M$; Number of preference pairs per input $K$; Number of iterations $T$; Hyperparameter $\lambda$.
for $t=0$ to $T-1$ do
       Initialize preference dataset $\mathcal{D}_{t}=\emptyset$;
       foreach $x_{i}\in\mathcal{D}$ do
             Initialize chosen code set $Y_{i}^{(c)}=\emptyset$;
             Initialize rejected code set $Y_{i}^{(r)}=\emptyset$;
            for $k=1$ to $M$ do
                   Generate response $\hat{y}_{i}^{k}\sim\pi_{t}(x_{i})$;
                   if Execution result of $\hat{y}_{i}^{k}$ matches ground truth $y_{i}$ then
                         Add $\hat{y}_{i}^{k}$ to $Y_{i}^{(c)}$;
                  else
                         Add $\hat{y}_{i}^{k}$ to $Y_{i}^{(r)}$;
            for $k=1$ to $K$ do
                   Randomly sample $\hat{y}_{i}^{c_{k}}$ from $Y_{i}^{(c)}$ (with replacement if $|Y_{i}^{(c)}|<K$);
                   Randomly sample $\hat{y}_{i}^{r_{k}}$ from $Y_{i}^{(r)}$ (with replacement if $|Y_{i}^{(r)}|<K$);
                   Add $(x_{i},\hat{y}_{i}^{c_{k}},\hat{y}_{i}^{r_{k}})$ to $\mathcal{D}_{t}$;
      Update model $\pi_{t+1}$ by minimizing the combined loss:
Algorithm 1 CodeLutra
Appendix B More information
B.1 Pseudo code of CodeLutra
We summarize our algorithm in implementation in the Algorithm 1. The algorithm operates iteratively, generating code responses for each input and leveraging both correct and incorrect code samples to construct a preference dataset. These comparisons serve as the basis for model refinement, where the model updates its parameters in each iteration. The process ensures that the model learns not only from correct code solutions but also from common mistakes, thereby improving its ability to generate high-quality code across diverse tasks. This iterative refinement, guided by self-generated comparative data, distinguishes CodeLutra from traditional supervised fine-tuning approaches.
B.2 Experimental setup
@@ B.3 Datasets
B.4 Examples for different datasets.
@@ B.5 Qualitative example
B.6 Setup of the baseline
@@
Figure 5: An example from the data query dataset from the BIRD (Wang et al., 2023a).
@@
Appendix C More results
C.1 Assessment ability
@@
C.2 Comparison of execution-based methods
To examine the limitations of previous methods on small LLMs, we evaluated SFT, SFT with self-debugging Chen et al. (2024), and CodeLutra on the Spider benchmark using Llama-3-8B as the base model. The results are summarized in Table 6.
@@ C.2 Comparison of execution-based methods
SELECT COUNT(CustomerID) FROM yearmonth WHERE Consumption > 1000
AND CustomerID IN (
    SELECT CustomerID FROM customers WHERE Currency = ’EUR’ );
SELECT COUNT(CustomerID) AS "Number of customers" FROM  "yearmonth"
WHERE  Consumption > 1000 AND     Currency = ’EUR’   GROUP BY CustomerID
C.3 Comparision with GPT-4 preference on code
To assess the reliability of using execution results versus GPT-4 preferences for constructing pairwise preferences, we conducted experiments on DS-1000 with Llama-3-8B as the base model. Specifically, we used GPT-4 to judge which answer in each pair is better, similar to the approach employed by RLAIF Lee et al. (2023). The results are presented in Table 7.
@@ C.3 Comparision with GPT-4 preference on code
Appendix D Limitation and future work
While CodeLutra enhances code generation performance by leveraging self-generated data, it exhibits limitations that warrant consideration. The framework focuses on the correctness of the generated code, overlooking other vital aspects such as efficiency, readability, and adherence to specific formal specifications, which are essential for practical applications. Additionally, CodeLutra treats all failed code attempts uniformly, without distinguishing between different types or severities of errors, potentially limiting the model’s ability to learn from more informative mistakes. To address the aforementioned limitations, future research related to CodeLutra should explore several key directions. Expanding the preference-guided refinement mechanism to incorporate additional criteria such as code efficiency and compliance with formal specifications would enhance the overall quality and utility of the generated code. Developing a more nuanced approach to categorizing and prioritizing failed code attempts based on the type and severity of errors could enable more targeted and effective learning, thereby improving the model’s ability to avoid similar mistakes in future generations. Exploring alternative evaluation methods, such as static code analysis or formal verification tools, could reduce the framework’s reliance on execution results and broaden its applicability to a wider range of tasks.

### v3

@@ 5.2 Main Results
We compare CodeLutra with baselines in code generation for the data query task, as shown in Table 1. We found that existing open-source LLMs like Llama-3-8B still have a significant performance gap in code generation for data queries compared to closed-source LLMs like GPT-4. Although supervised fine-tuning can help bridge this gap—e.g., SFT increases the EX of Llama-3-8B on Spider from 59.3% to 67.9%—there remains a notable difference with GPT-4’s 74.4%. Through our refinement framework, Llama-3-8B after four iterations exceeded SFT performance by 16.9% and even outperforms GPT-4 with an execution accuracy of 76.6%. Additionally, on the more challenging BIRD dataset, after three iterations, CodeLutra significantly improved the EX of the base model from 22.3 to 43.1, achieving performance very close to GPT-4. The similar performance observed between iterations 3 and 4 can be attributed to model convergence, as the model approaches its optimal state with no further gains in later iterations.
@@ 5.3 More experiments
Our framework CodeLutra leverages both positive and negative answer pairs to iteratively improve model performance, particularly by minimizing the generation of incorrect responses. But what happens when we omit the negative samples and rely solely on supervised fine-tuning using positive samples generated by the model? In this ablation, we compare the performance of our objective 4 with a model trained with $\mathcal{L}_{\text{SFT}}(\pi;\mathcal{D}_{t}^{c})$. As seen in Figure 3(a), without negative samples (purple line), the model’s performance plateaus across iterations, remaining close to the baseline. In contrast, incorporating negative samples (blue line) leads to steady performance improvements over successive iterations. This ablation confirms that including negative samples is critical to refining the model’s ability to distinguish between optimal and suboptimal responses, significantly boosting overall performance.
@@
Table 2: Ablations on the SFT on the correct answers.
Methods Spider DS1000
$\mathcal{L}_{\text{DPO}}$ 17.2 12.4
Ours 76.6 48.6
@@ 6 Further Analysis on CodeLutra
SELECT COUNT(CustomerID) FROM yearmonth WHERE Consumption > 1000
AND CustomerID IN (
    SELECT CustomerID FROM customers WHERE Currency = ’EUR’ );
SELECT COUNT(CustomerID) AS "Number of customers" FROM  "yearmonth"
WHERE  Consumption > 1000 AND     Currency = ’EUR’   GROUP BY CustomerID
@@
Based on the responses of the initial model $\pi_{0}$, we divide the test set into a correct set and an error set. We track the quality trends of the model in these two sets across iterations. Using the cosine similarity metric based on the BLEURT embedding proposed by (Sellam et al., 2020), we calculate similarities denoted as $\text{sim}{\pi_{t}}(\hat{y},y_{\text{gt}})$ for the model fine-tuned over $t$ iterations. Here $\hat{y}$ denotes the model generation, and $y_{\text{gt}}$ is the ground truth solution. As shown in Figure 4(b), we observe that the similarity between the model’s output on the correct set and the ground truth remains stable (see purple bars), while with each iteration, the similarity between the model’s output on the error set and the ground truth increases significantly—from 0.48 to 0.54. This indicates that CodeLutra helps the base model improve outputs on error set, while the outputs on correct cases remain qualitatively stable.
@@ 7 Conclusion
Appendix A More information
A.1 Experimental setup
@@ A.2 Datasets
A.3 Examples for different datasets
@@ A.4 Qualitative example
A.5 Model checkpoints and reproducibility
To ensure reproducibility of our experiments, Table 5 provides the exact model checkpoints used for each model in our evaluation. For proprietary models, we include the specific version and access method.
Table 5: Exact model checkpoints used in our experiments.
Model Checkpoint/Version
Llama-3-8B meta-llama/Llama-3-8B-Instruct
Llama-3-70B meta-llama/Llama-3-70B-Instruct
CodeLlama-7B codellama/CodeLlama-7b-Instruct-hf
StarCoder-15B bigcode/starcoder
Gemma-7B google/gemma-7b-it
Codex code-davinci-002 (via Azure OpenAI Service)
GPT-3.5 gpt-3.5-turbo-0613
GPT-4 gpt-4-0613
A.6 Setup of the baseline
@@
Figure 6: An example from the data query dataset from the BIRD (Li et al., 2024).
@@
Appendix B More results
B.1 Assessment ability
@@
B.2 Comparison of execution-based methods
To examine the limitations of previous methods on small LLMs, we evaluated SFT, SFT with self-debugging (Chen et al., 2024), and CodeLutra on the Spider benchmark using Llama-3-8B as the base model. The results are summarized in Table 7.
@@ B.2 Comparison of execution-based methods
B.3 Comparision with GPT-4 preference on code
To assess the reliability of using execution results versus GPT-4 preferences for constructing pairwise preferences, we conducted experiments on DS-1000 with Llama-3-8B as the base model. Specifically, we used GPT-4 to judge which answer in each pair is better, similar to the approach employed by RLAIF (Lee et al., 2023). The results are presented in Table 8.
@@ B.3 Comparision with GPT-4 preference on code
B.4 Additional Benchmark Results
To demonstrate the generalizability of our approach beyond the primary datasets used in our main experiments, we evaluated CodeLutra on the MBPP benchmark from EvalPlus (Liu et al., 2023a). Using Llama-3-8B as our base model, we compared the performance of the base model, standard supervised fine-tuning (SFT), and our CodeLutra approach.
Methods Accuracy (%)
Base 52.3
SFT 53.1
CodeLutra 55.4
Table 9: Performance comparison on the MBPP benchmark from EvalPlus using Llama-3-8B.
These results confirm that CodeLutra’s effectiveness extends beyond our primary datasets, demonstrating its broad applicability across different code generation tasks and benchmarks.
Appendix C Limitation and future work
While CodeLutra enhances code generation performance by leveraging self-generated data, it exhibits limitations that warrant consideration. The framework focuses on the correctness of the generated code, overlooking other vital aspects such as efficiency, readability, and adherence to specific formal specifications, which are essential for practical applications. Additionally, CodeLutra treats all failed code attempts uniformly, without distinguishing between different types or severities of errors, potentially limiting the model’s ability to learn from more informative mistakes. Furthermore, our current sampling strategy for preference pairs could be refined—while we sample with replacement when there are fewer than K examples, statistical theory suggests that sampling with replacement even when more examples are available may be beneficial, as it would better approximate sampling from the underlying distribution (similar to bootstrapping in statistics). To address the aforementioned limitations, future research related to CodeLutra should explore several key directions. Expanding the preference-guided refinement mechanism to incorporate additional criteria such as code efficiency and compliance with formal specifications would enhance the overall quality and utility of the generated code. Developing a more nuanced approach to categorizing and prioritizing failed code attempts based on the type and severity of errors could enable more targeted and effective learning, thereby improving the model’s ability to avoid similar mistakes in future generations. Exploring alternative evaluation methods, such as static code analysis or formal verification tools, could reduce the framework’s reliance on execution results and broaden its applicability to a wider range of tasks. Additionally, investigating optimal sampling strategies for preference pairs, including consistent sampling with replacement regardless of sample size, may further improve the statistical properties of our approach.